OpenStudy (anonymous):

Solve for x: −3|x + 7| = −12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (anonymous):

I think it is C. x = -3, x = 11

OpenStudy (whovianchick):

Divide both sides by -3, then find x+7=4, and x+7=-4 X+7=4 x+7=-4 x+7-7=4-7 x+7-7=-4-7 x=-3 x=-11 B. x=-3, x=-11

OpenStudy (anonymous):

Really you

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

ok hopefully i can trust this answer

OpenStudy (anonymous):

@whovianchick

OpenStudy (whovianchick):

Haha yes you can −3|x + 7| = −12 −3|x + 7|/-3 = −12/-3 |x + 7| = 4 X+7=4 x+7=-4 x+7-7=4-7 x+7-7=-4-7 x=-3 x=-11 B. x=-3, x=-1 There's all the work

OpenStudy (whovianchick):

And, plug both the answers in. They work. Aaaaaaaalways plug the solutions in to check for extraneous solutions.

OpenStudy (anonymous):

x=-3 or x=-11

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

:P